# Perturbation-based Effect Measures for Compositional Data

### v3

@@ 2.2 Perturbation effects on the simplex
To make the discussion more concrete, we now provide several explicit perturbation effects on the simplex using the perturbations constructed in Section 2.1. We motivate each perturbation effect with an example application to showcase potential use cases. A list summarizing all proposed perturbation effects together with their respective derivative-isolating reparametrizations that are used for efficient estimation is provided in Table 2. We encourage practitioners to develop further application-specific perturbations that optimally target the research question they are interested in.
@@ 2.2.1 Effects of changes in individual components
Intuitively, this stems from the fact that knowing $Z^{-j}$ implies that we also know $Z^{j}$ via the relation $Z^{j}=1-\|Z^{-j}\|_{1}$ hence $Z^{j}$ can never reveal any additional information regarding $Y$. We therefore propose – for a fixed component $j$ – to consider perturbations that increase the $j$th component while decreasing all remaining components proportionally. We define the compositional feature influence for the $j$th feature or simply $\mathrm{CFI}^{j}$ as $\tau_{\psi}$ for any directional perturbation with endpoint $e_{j}$. To that end, we already considered these perturbations in Example 3 where we proposed two possible speeds leading to slightly different interpretations of the average perturbation effects: $\mathrm{CFI}^{j}_{\mathrm{unit}}$ is the perturbation that moves at a constant (unit) speed towards the vertex and $\mathrm{CFI}^{j}_{\mathrm{mult}}$ is the perturbation that moves at what we call multiplicative speed (see Example 3 for more intuition on these speeds). $\mathrm{CFI}_{\mathrm{mult}}^{j}$ has previously been proposed by Huang et al. (2023), however without an efficient estimation procedure. Huang et al. (2023) also prove that when a log-contrast regression model is well-specified the $j$th coefficient in the model is equal to $\mathrm{CFI}_{\mathrm{mult}}^{j}$ (this proof is also included in Section S6 of the supplementary material). A crucial difference between unit speed and multiplicative speed perturbations is how they account for observations with $Z^{j}=0$. As the multiplicative speed is 00 in this case such points do not contribute to $\mathrm{CFI}^{j}_{\mathrm{mult}}$, however, they do contribute to $\mathrm{CFI}^{j}_{\mathrm{unit}}$.
@@ 2.2.2 Effects of changes in diversity
As a second example we consider perturbation effects that capture how a response is associated with increases in diversity. Such effects are often relevant in social sciences and econometrics, e.g., when considering racial or income distributions. As we outlined in the introduction, it can be misleading to only model summary statistics of diversity and ignore the remaining compositional structure as this structure can confound the relationship between the diversity measure and the response. Instead, we propose to consider the effect of directional perturbations that push towards $z_{\mathrm{cen}}$ (the centering perturbations defined in Section 2.1.1) and call $\tau_{\psi}$ for such perturbations a compositional diversity influence ($\mathrm{CDI}$). We denote the effect by $\mathrm{CDI}_{\mathrm{unit}}$ if the perturbation has unit speed and by $\mathrm{CDI}_{\mathrm{Gini}}$ if the perturbation was constructed as in Example 5 such that $L$ is equal to the negative Gini coefficient. The construction in Example 5 can be adapted to many popular diversity measures and hence allows practitioners to use their favorite diversity summary statistics. In Section 4.1.1, we create a semi-synthetic dataset using data from the $1994$ US census to analyze the effect of racial diversity on income and show how $\mathrm{CDI}_{\mathrm{Gini}}$ avoids the confounding caused by naively including the Gini coefficient in a regression. In Section 4.1.2 we analyze the effect of racial diversity on grades in New York schools to further illustrate the use of $\mathrm{CDI}_{\mathrm{Gini}}$.
@@
2.2.3 Effects of amalgamations
As a third example we consider perturbation effects that capture the influence of general amalgamations. As an example, consider a study where data is collected on how individuals spend their time each day (e.g., hours spent sleeping, eating, working, doing sports, etc.) as well as a response variable (e.g., general happiness). Assume now that we want to use this data to investigate the relationship between time spent on sports and happiness. As above, we are again interested in the effect of an individual component, so we might consider estimating a $\mathrm{CFI}^{j}$. However, this would capture the effect of increasing time spent on sports while all other activities are down-scaled proportionally. For example, time spent sleeping and eating would both be decreased if time doing sports is increased which is not realistic. To address this, we can instead consider a general $(A\to B)$-amalgamating perturbation (as defined in Section 2.1.1), where $B=\{\text{sport}\}$ and $A$ corresponds to activities which are reduced to make increased time for sports. Perturbing in this fashion keeps the components in $[d]\setminus(A\cup B)$ fixed (e.g., eating and drinking) while down-scaling the activities in $A$. We call directional perturbation effects of this form compositional amalgamation influences, or $\mathrm{CAI}^{A\to B}$ for short. As with the $\mathrm{CFI}^{j}$, we use subscripts to denote the speed of the perturbation, so we have $\mathrm{CAI}_{\mathrm{unit}}^{A\to B}$ and $\mathrm{CAI}^{A\to B}_{\mathrm{mult}}$ (see Example 4).
Analogously, we could be interested in the effect of removing particular activities (e.g. gambling) entirely and spending the time on desirable activities (e.g. sports) while fixing time spent on a third group of activities (e.g. eating, sleeping). We can capture this effect using a binary amalgamating perturbation $\psi$. We call the resulting average binary perturbation effect for $\psi$ an $(A\to B)$-compositional amalgamation effect, or $\mathrm{CAE}^{A\to B}$ for short.
@@ 2.2.3 Effects of amalgamations
$\mathrm{CDI}_{\mathrm{Gini}}$ see (10) $z_{\mathrm{cen}}$ $-G(z)$, see (9)
$\mathrm{CAE}^{A\to B}$ binary $\mathsf{A}_{A\to B}(z)$ $\mathds{1}_{\{z^{A}=0\}}$
@@
2.3 Perturbations as causal quantities
Perturbations are a priori non-causal quantities and should not be confused with (causal) interventions. To make this distinction clear, we now discuss how our perturbation framework can be used for causal inference. Causal models (Pearl, 2009; Rubin, 2005) provide a mathematical framework for modeling changes to the data-generating mechanism – often called interventions. In contrast to an (observational) statistical model, which assumes a single observational distribution that generated the data, a causal model additionally postulates the existence of a collection of interventional distributions that describe the data-generating mechanism under hypothetical interventions. The power of causal models is that they allow specification of rigorous mathematical conditions under which (parts of) the interventional distributions can be identified (or computed) from the observational distribution. Unlike interventions, perturbations describe hypothetical changes to the predictor space (in this work $\Delta^{d-1}$, but the same ideas hold more generally) and not to the data-generating mechanism. To make this more precise and show how to use perturbations for causal inference, we use the structural causal model (SCM) framework (Pearl, 2009; Bongers et al., 2021). Similar considerations, however, also apply with the potential outcome framework.
As an example, consider the SCM over the variables $(H,Z,Y)\in\mathbb{R}^{q}\times\Delta^{d-1}\times\mathbb{R}$ given by
Assume now we observe data generated by $\mathcal{M}$ and are interested in understanding how $Z$ causally affects $Y$. One target parameter might be to consider the expected value of $Y$ under an intervention that sets $Z$ to a fixed value, using do-notation (Pearl, 2009), this can be expressed as $\mathbb{E}[Y\mid\operatorname{do}(Z=z)]=\mathbb{E}[g(z,H)]$. For larger $d$ the full causal relation might be too ambitious of a goal, and it may be more tractable to look at the effect of individual components. Since the vector $Z$ is compositional, one needs to be careful when expressing summary interventions. Perturbations provide a solution to this as they allow us to express sensible ways of changing the predictors. More specifically, let $\psi$ be a perturbation as defined in Definition 1. Then, for all $\gamma$ sufficiently small, consider the intervention $\operatorname{do}(Z=\psi(Z,\gamma))$, which in $\mathcal{M}$ replaces the $Z$-structural assignment by $Z\leftarrow\psi(g(H,\varepsilon_{Z}),\gamma)$. In a similar way as to how we previously used the conditional expectation of $Y$ under the perturbation, we can now use the expected value of $Y$ under the perturbation-induced intervention, i.e., $\mathbb{E}[Y\mid\operatorname{do}(Z=\psi(Z,\gamma))]$. Here, we only consider interventions on the entire compositional vector $Z$, which means that it is not possible to model how causal effects propagate within the composition itself. The corresponding causal versions of the target parameters in Definition 2 are then the causal average binary perturbation effect
and the causal average directional perturbation effect
$\mathbb{E}[Y\mid\operatorname{do}(Z=\psi(Z,\gamma))]=\mathbb{E}\big{[}\mathbb{E}[Y\mid Z=\psi(Z,\gamma),H]\big{]},$ (11)
and hence is identifiable. The formula in (11) is a special case of the adjustment formula (e.g., Pearl, 2009), but other ways of identifying this effect also exist (e.g., propensity score matching). It is easy to adapt the estimators proposed in this work to such settings by conditioning on additional confounding covariates $H$. In particular, starting from the reparametrization into $(L,W)$ one only needs to add additional observed confounders alongside $W$. This is for example done in the numerical experiments in Section 4.1.
Finally, let us shortly discuss under which assumptions the causal and non-causal average perturbation effects coincide. Obviously, one sufficient condition is to assume that
$\mathbb{E}[Y\mid\operatorname{do}(Z=\psi(Z,\gamma))]=\mathbb{E}[Y\mid Z=\psi(Z,\gamma)].$ (12)
Such a condition holds by definition whenever $Z$ is an exogenous variable (i.e., does not depend on other variables in its structural assignment) but can also hold more generally. A common assumption to ensure (12) is to assume that $Z$ and $Y$ are not confounded, that is, $Z$ and $Y$ have no common causes (this is known as ignorability in the potential outcome framework). In the SCM $\mathcal{M}$ this would for example mean that $H$ does not exist or either $g$ or $h$ do not depend on $H$. In many cases, this is not satisfied and one should instead carefully argue for which causal assumptions are reasonable and adjust the estimation accordingly, e.g., by conditioning on additional covariates as in (11).
@@ 3 Estimating perturbation effects
In this section, we describe how to estimate and test hypotheses about the target parameters $\tau_{\psi}$ and $\lambda_{\psi}$. We will not explicitly restrict ourselves to any one perturbation $\psi$ as the approaches described are agnostic to this choice. However, we will assume that any directional perturbation comes with a derivative-isolating reparametrization $\phi=(\phi^{L},\phi^{W})$ and for any binary perturbation we define $\phi^{L}:z\mapsto\mathds{1}_{\{z=\psi(z,1)\}}$ and $\phi^{W}:z\mapsto\psi(z,1)$. This allows us to reparametrize our estimation problem from a compositional data problem with $n$ i.i.d. observations of $(Y,Z)\in\mathbb{R}\times\Delta^{d-1}$ into a generic problem with $n$ i.i.d. observations of $(Y,L,W)$ with $W\in\mathcal{W}$ and $L\in\mathbb{R}$, where our interest lies in the relationship between $Y$ and $L$ when controlling for $W$. We only present an overview of the different estimators in this section and postpone a detailed description of the algorithms and theory to Section 5.
@@ 3.1 Efficient estimation using semiparametric theory
We now give a brief introduction to semiparametric estimation of statistical functionals and, in particular, we describe how naive plug-in estimation of target parameters based on nonparametric methods rarely leads to efficient tests and how to remedy this. This perspective is not new and goes back to at least Pfanzagl (1982) with several more recent monographs and reviews (Bickel et al., 1993; Tsiatis, 2006; Chernozhukov et al., 2018; Kennedy, 2023). The recent surge in the popularity of these methods is fueled by the advent of powerful machine learning methods that use flexible black box models to learn the nuisance functions involved in the estimation of functionals of interest. For illustration, we focus on the estimation of $\tau_{\psi}$, that is, the estimation of
@@
Figure 3: Plot of $95\%$ confidence intervals for a plug-in, one-step and cross-fit one-step estimator of $\lambda_{\psi}$ across 100 repetitions in the partially linear, binary simulation setting described in Section S4.2 of the supplementary material with $d=15$ and $n=1000$. The black dashed line indicates the true value of the parameter and we highlight the confidence intervals that do not contain the true parameter. The confidence intervals are ordered first by whether the true value is in the interval, then by the value of the estimate.
@@ 3.2 Efficient estimation of perturbation effects
In this section, we present our proposals for estimating perturbation effects. We postpone the theoretical analysis and a detailed description of cross-fitting to Section 5 and only provide a high-level overview here. All the presented estimators involve regressions that can be performed by arbitrary (sufficiently predictive) regression methods.
@@ 3.2.1 Nonparametric estimation of the binary perturbation effect
A full algorithm is given in Algorithm 2. It is worth noting that this estimator is sensitive to the estimation of the propensity score $\pi$ which can be difficult when the dimension of $W$ is large or the relationship between $L$ and $W$ is complicated (see Section S3.1 of the supplementary material for an illustration of this).
@@ 3.2.2 Nonparametric estimation of the directional perturbation effect
where $\widehat{f}$ is a differentiable estimator of $f:(\ell,w)\mapsto\mathbb{E}[Y\,|\,L=\ell,W=w]$ and $\widehat{\rho}$ is an estimate of the conditional score function $\rho$ of $L$ given $W$, that is, an estimate of the function $\rho:(\ell,w)\mapsto\partial_{\ell}\log(p(\ell|w))$ where $p$ denotes the conditional density of $L$ given $W$. A full algorithm for the estimation is given in Algorithm 1 below. Both the estimation of $f$ with computable derivative $\partial_{\ell}f$ and $\rho$ can be challenging and we show in Section 4 that the overall estimator (15) is sensitive to both estimates (see Section S3.1 of the supplementary material for an illustration of this).
For the estimation of $f$ we require that the estimator $\widehat{f}$ is differentiable and with computable derivatives (which additionally need to approximate the true derivative at a sufficiently fast rate). This excludes the direct use of many popular machine learning methods such as random forests or boosted trees. One solution is to start from a potentially non-differentiable pilot estimate of $f$ and smooth it to get a differentiable estimate for which derivatives can be computed. Such a procedure based on kernel-smoothing has been proposed by Klyne and Shah (2023). Here, we use an alternative smoothing method which uses random forests and local polynomials. A full description of our smoothing method is provided in Section S3.3 of the supplementary material. Finally, for the estimation of $\rho$, we follow the proposal in Klyne and Shah (2023) and assume a location-scale model. This requires the estimation of the conditional mean and variance of $L$ given $W$ (referred to as $\eta$ in Algorithm 1). Further details on the location-scale based estimate of $\rho$ is given in Section S3.4 of the supplementary material.
@@ 3.2.3 Partially linear estimation of perturbation effects
While the estimators in the previous sections make only mild assumption on the data-generating model, they rely on accurate estimates of quantities that may be difficult to estimate, such as derivatives of densities and regression functions. One way to avoid these complicated estimations is by additionally assuming a partially linear model. For both binary and directional perturbations (with a derivative-isolating reparametrization), we can assume a partially linear model on $L$ and $W$, that is, the existence of $h$ and $\theta$ such that
@@
where $\widehat{g}$ is an estimator of $g:w\mapsto\mathbb{E}[Y\,|\,W=w]$ and $\widehat{m}$ is an estimator of $m:w\mapsto\mathbb{E}[L\,|\,W=w]$. A full algorithm is given in Algorithm 3. In contrast to the estimators from the previous sections, this estimator only relies on estimating two conditional mean functions. This comes at the expense of assuming that there is no interaction effect between $L$ and $W$ on $Y$ and that the effect of $L$ is linear on $Y$. Even when these assumptions are not satisfied, the estimator may still provide a reasonable estimate of the effect of $L$ on $W$ and we found in our experiments that the estimator is more robust to errors in the estimation of the nuisance functions than the estimators above. In fact, one can show that the quantity $\mathbb{E}[(\mathbb{E}[Y\,|\,L,W]-\theta L-h(W))^{2}]$ is minimized by the population version of the partially linear model estimate, so one can think of the partially linear model as a ‘best approximation’ to a generic nonparametric model (see Proposition S1 in Section S3.1.3 of the supplementary material for a proof of this fact).
@@ 4 Numerical experiments
We now present numerical experiments illustrating the usage and performance of our methods. In Section 4.1.1 we analyze the effect of racial diversity on income on a semi-synthetic dataset based on data from the 1994 US census. In Section 4.1.2, we analyze the effect of racial diversity on grades in New York schools. In Section 4.2 we investigate variable influence (importance) measures for the prediction of BMI from gut microbiome data. In addition to the experiments included here, Section S4.2 of the supplementary material contains an investigation of the performance of the semiparametric estimators discussed in Section 3 when the conditioning variable $W$ is compositional. The code for all of our experiments is available at https://github.com/ARLundborg/PerturbationCoDa. For all regressions we use the scikit-learn package (Pedregosa et al., 2011) except for the $\ell^{1}$-penalized log-contrast regression in Section 4.2 where we used the c-lasso package (Simpson et al., 2021).

### v4

@@ 2.2 Perturbation effects on the simplex
To make our proposal more concrete, we now provide several explicit perturbation effects on the simplex using the perturbations constructed in Section 2.1. We motivate each perturbation effect with an example application to showcase potential use cases. In Section S2.6 we describe an additional class of perturbation effects based on general amalgamating perturbations.
@@ 2.2.1 Effects of changes in individual components
Intuitively, this stems from the fact that knowing $Z^{-j}$ implies that we also know $Z^{j}$ via the relation $Z^{j}=1-\|Z^{-j}\|_{1}$ hence $Z^{j}$ can never reveal any additional information regarding $Y$. We therefore propose – for a fixed component $j$ – to consider perturbations that increase the $j$th component while decreasing all remaining components proportionally. We define the compositional feature influence for the $j$th feature or simply $\mathrm{CFI}^{j}$ as $\tau_{\psi}$ for any directional perturbation with endpoint $e_{j}$. We already considered these perturbations in Example 1 where we examined two possible speeds leading to slightly different interpretations of the average perturbation effects: $\mathrm{CFI}^{j}_{\mathrm{unit}}$ is the perturbation that moves at a constant (unit) speed towards the vertex and $\mathrm{CFI}^{j}_{\mathrm{mult}}$ is the perturbation that moves at what we call multiplicative speed. $\mathrm{CFI}_{\mathrm{mult}}^{j}$ has previously been proposed by Huang et al. (2023), however without an efficient estimation procedure. Huang et al. (2023) also prove that when a log-contrast regression model is well-specified the $j$th coefficient in the model is equal to $\mathrm{CFI}_{\mathrm{mult}}^{j}$ (this proof is also included in Section S2.8). A crucial difference between unit speed and multiplicative speed perturbations is how they account for observations with $Z^{j}=0$. As the multiplicative speed is 00 in this case such points do not contribute to $\mathrm{CFI}^{j}_{\mathrm{mult}}$, however, they do contribute to $\mathrm{CFI}^{j}_{\mathrm{unit}}$.
@@ 2.2.2 Effects of changes in diversity
As a second example we consider perturbation effects that capture how a response is associated with increases in diversity. Such effects are often relevant in social sciences and econometrics, e.g., when considering racial or income distributions. As we outlined in the introduction, it can be misleading to only model summary statistics of diversity and ignore the remaining compositional structure as this structure can confound the relationship between the diversity measure and the response. Instead, we propose to consider the effect of directional perturbations that push towards $z_{\mathrm{cen}}$ (centering perturbations as defined in Section 2.1.1) and call $\tau_{\psi}$ for such perturbations a compositional diversity influence ($\mathrm{CDI}$). We denote the effect by $\mathrm{CDI}_{\mathrm{unit}}$ if the perturbation has unit speed and by $\mathrm{CDI}_{\mathrm{Gini}}$ if the perturbation was constructed as in Example 2 such that $L$ is equal to one minus the Gini coefficient. The construction in Example 2 can be adapted to many popular diversity measures and hence allows practitioners to use their favorite diversity summary statistics. In Section 4.1.1, we create a semi-synthetic dataset using data from the $1994$ US census to analyze the effect of racial diversity on income and show how $\mathrm{CDI}_{\mathrm{Gini}}$ avoids the confounding caused by naively including the Gini coefficient in a regression. In Section 4.1.2 we analyze the effect of racial diversity on grades in New York schools to further illustrate the use of $\mathrm{CDI}_{\mathrm{Gini}}$.
@@
$\mathrm{CDI}_{\mathrm{Gini}}$ see (8) $z_{\mathrm{cen}}$ $1-G(z)$, see (7)
@@ 3 Estimating perturbation effects
In this section, we describe how to estimate and construct confidence intervals for the target parameters $\tau_{\psi}$ and $\lambda_{\psi}$. We will not explicitly restrict ourselves to any one perturbation $\psi$ as the approaches described are agnostic to this choice. However, we will assume that any directional perturbation comes with a derivative-isolating reparametrization $\phi=(\phi^{L},\phi^{W})$ and for any binary perturbation we define $\phi^{L}:z\mapsto\mathds{1}_{\{z=\psi(z,1)\}}$ and $\phi^{W}:z\mapsto\psi(z,1)$. This allows us to reparametrize our estimation problem from a compositional data problem with $n$ i.i.d. observations of $(Y,Z)\in\mathbb{R}\times\Delta^{d-1}$ into a generic problem with $n$ i.i.d. observations of $(Y,L,W)$ with $W\in\mathcal{W}$ and $L\in\mathbb{R}$, where our interest lies in the relationship between $Y$ and $L$ when controlling for $W$. We present an overview of the different estimators in this section and provide further details of the algorithms and theory in Section S4.1.
@@ 3.1 Efficient estimation using semiparametric theory
We now give a brief introduction to semiparametric estimation of statistical functionals and, in particular, we describe how naive plug-in estimation of target parameters based on nonparametric methods rarely leads to efficient confidence intervals and how to remedy this. This perspective is not new and goes back to at least Pfanzagl (1982) with several more recent monographs and reviews (Bickel et al., 1993; Tsiatis, 2006; Chernozhukov et al., 2018; Kennedy, 2023). The recent surge in the popularity of these methods is fueled by the advent of powerful machine learning methods that use flexible black box models to learn the nuisance functions involved in the estimation of functionals of interest. For illustration, we focus on the estimation of $\tau_{\psi}$, that is, the estimation of
@@
Figure 1: Plot of $95\%$ confidence intervals for a plug-in, one-step and cross-fit one-step estimator of $\lambda_{\psi}$ across 100 repetitions in the partially linear, binary simulation setting described in Section S5.2 with $d=15$ and $n=1000$. The black dashed line indicates the true value of the parameter and we highlight the confidence intervals that do not contain the true parameter. The confidence intervals are ordered first by whether the true value is in the interval, then by the value of the estimate.
@@ 3.2 Efficient estimation of perturbation effects
In this section, we present our proposals for estimating perturbation effects. Further details and full algorithms can be found in Section S4.1. All the presented estimators involve regressions that can be performed by arbitrary (sufficiently predictive) regression methods.
@@ 3.2.2 Nonparametric estimation of the directional perturbation effect
where $\widehat{f}$ is a differentiable estimator of $f:(\ell,w)\mapsto\mathbb{E}[Y\,|\,L=\ell,W=w]$ and $\widehat{\rho}$ is an estimate of the conditional score function $\rho$ of $L$ given $W$, that is, an estimate of the function $\rho:(\ell,w)\mapsto\partial_{\ell}\log(p(\ell|w))$ where $p$ denotes the conditional density of $L$ given $W$.
For the estimation of $f$ we require that the estimator $\widehat{f}$ is differentiable and with computable derivatives (which additionally need to approximate the true derivative at a sufficiently fast rate). This excludes the direct use of many popular machine learning methods such as random forests or boosted trees. One solution is to start from a potentially non-differentiable pilot estimate of $f$ and smooth it to get a differentiable estimate for which derivatives can be computed. Such a procedure based on kernel-smoothing has been proposed by Klyne and Shah (2023). Here, we use an alternative smoothing method which uses random forests and local polynomials. Finally, for the estimation of $\rho$, we follow the proposal in Klyne and Shah (2023) and assume a location-scale model. This requires the estimation of the conditional mean and variance of $L$ given $W$ (referred to as $\eta$ in Algorithm S1). Further details on our smoothing method and the location-scale based estimate of $\rho$ are given in Sections S4.5 and S4.6, respectively.
@@ 3.2.3 Partially linear estimation of perturbation effects
While the estimators in the previous sections make only mild assumption on the data-generating model, they rely on accurate estimates of quantities that may be difficult to estimate, such as derivatives of densities and regression functions (see Section S4.3 for an illustration of this sensitivity). One way to avoid these complicated estimations is by additionally assuming a partially linear model. For both binary and directional perturbations (with a derivative-isolating reparametrization), we can assume a partially linear model on $L$ and $W$, that is, the existence of $h$ and $\theta$ such that
@@
where $\widehat{g}$ is an estimator of $g:w\mapsto\mathbb{E}[Y\,|\,W=w]$ and $\widehat{m}$ is an estimator of $m:w\mapsto\mathbb{E}[L\,|\,W=w]$. In contrast to the estimators from the previous sections, this estimator only relies on estimating two conditional mean functions. This comes at the expense of assuming that there is no interaction effect between $L$ and $W$ on $Y$ and that the effect of $L$ is linear on $Y$. Even when these assumptions are not satisfied, the estimator may still provide a reasonable estimate of the effect of $L$ on $W$ and we found in our experiments that the estimator is more robust to errors in the estimation of the nuisance functions than the estimators above. In fact, one can show that the quantity $\mathbb{E}[(\mathbb{E}[Y\,|\,L,W]-\theta L-h(W))^{2}]$ is minimized by the population version of the partially linear model estimate, so one can think of the partially linear model as a ‘best approximation’ to a generic nonparametric model (see Proposition S4 in Section S4.3.3 for a proof).
@@ 4 Numerical experiments
We now present numerical experiments illustrating the usage and performance of our methods. In addition to the experiments included here, Section S5.2 contains an investigation of the performance of the semiparametric estimators discussed in Section 3 when the conditioning variable $W$ is compositional. The code for all of our experiments is available at https://github.com/ARLundborg/PerturbationCoDa. For all regressions we use the scikit-learn package (Pedregosa et al., 2011) except for the $\ell^{1}$-penalized log-contrast regression in Section 4.2 where we used the c-lasso package (Simpson et al., 2021).